Circular No. 6494

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/1996 U1 (SHOEMAKER-LEVY 3) G. Hahn, Institut für Planetenerkundung, DLR Forschungszentrum,

Berlin, reports the recovery of comet P/1991 C1 (= 1990 XXVII = 1991e) by A. Maury, M. Lundström and himself in the course of the OCA–DLR Asteroid Survey using the 0.9-m Schmidt telescope + CCD at Caussols. The comet was diffuse, with no tail, in moderate to poor seeing. The nominal correction (ignoring perturbations between now and the epoch) to the prediction by S. Nakano on MPC 25183 is ΔT = −0.1 day.

1996 UT  α_{2000}  δ_{2000}  m_{1}
Oct. 17.99750  1^{h}39^{m}09.^{s}25  +17^{o}39′57′′.9  19.3
18.00267  1 39 08.97  +17 39 56.9  19.2
18.01193  1 39 08.60  +17 39 54.9  19.3
18.92657  1 38 29.45  +17 36 34.7  19.3
18.93547  1 38 29.09  +17 36 33.2  19.2
18.94122  1 38 28.81  +17 36 32.1  19.2

The following improved orbital elements, by the undersigned, satisfy 25 observations 1991–1996 with mean residual 0′′.6:

Epoch = 1998 Mar. 8.0 TT

T = 1998 Mar. 4.9475 TT  ω = 181.^{o}3886
e = 0.247914  Ω = 303.7122 }2000.0
q = 2.817332 AU  i = 5.0086

a = 3.746025 AU n^{o} = 0.1359402 P = 7.250 years

1996TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Oct. 14 1^{h}41.^{m}98 +17^{o}53.^{′}9 2.649 3.633 168.^{o}6 3.^{o}1 19.2
24 1 34.86 +17 17.1 2.621 3.610 172.6  2.0 19.2
Nov. 3 1 27.96 +16 35.4 2.623 3.587 164.1  4.3 19.1
13 1 21.90 +15 52.8 2.654 3.565 153.1  7.2 19.1
23 1 17.24 +15 13.7 2.711 3.542 142.1  9.9 19.2
Dec. 3 1 14.33 +14 41.6 2.790 3.519 131.3 12.2 19.2
13 1 13.36 +14 19.0 2.888 3.497 120.8 14.0 19.2
23 1 14.38 +14 07.4 2.999 3.474 110.9 15.3 19.3

COMET C/1995 O1 (HALE-BOPP) Corrigendum. On IAUC 6493, line 6, for Q(NH_{2})/Q(H_{2}O) = 0.1 read

Q(NH_{2})/Q(H_{2}O) = 0.001.

1996 October 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden